# Negative cosmological constant in the dark energy sector: tests from JWST photometric and spectroscopic observations of high-redshift galaxies

### v1

@@ 2 Dark Energy models
Following the earlier work of Adil et al. (2023), we consider a DE sector consisting of a cosmological constant $\Lambda\gtrless 0$ which in principle can take either sign (with positive or negative signs corresponding to a dS or AdS vacuum respectively), and with associated energy density $\rho_{\Lambda}=\Lambda/8\pi G$ of the same sign. On top of this dS or AdS vacuum we place an evolving DE component with strictly positive energy density $\rho_{x}(z)>0$. Rather than committing to a specific microphysical model for the evolving DE component, we assume that this is described by a time-evolving EoS $w_{x}(z)$ of the Chevallier-Polarski-Linder (CPL) form (Chevallier & Polarski, 2001; Linder, 2003):
$w_{x}(z)=w_{0}+w_{a}\frac{z}{1+z}\,.$ (1)
There are various reasons why we adopt the widely used CPL parametrization, ranging from its highly manageable 2-dimensional nature, to its direct connection to several physical DE models (including several quintessence DE models, see e.g. Linder, 2003, 2006, 2008a, 2008b; Scherrer, 2015) and, last but not least, for ease of comparison to the earlier work of Adil et al. (2023) [^8] See Colgáin et al. (2021) for a recent discussion on potential shortcomings of the CPL parametrization. For completeness, we note that a number of other parametrizations for the EoS of dynamical DE components have been proposed in the literature, see e.g. Efstathiou (1999); Jassal et al. (2005); Gong & Zhang (2005); Barboza & Alcaniz (2008); Ma & Zhang (2011); Pantazis et al. (2016); Yang et al. (2018b); Pan et al. (2018); Yang et al. (2019); Singh et al. (2023) (see also Perkovic & Stefancic, 2020, for discussions on the theoretical viability of these parametrizations). We refer to the evolving component being quintessence-like (phantom) at a given epoch if, at the redshift in question, $w(z)>-1$ ($<-1$). While $w_{0}$ controls the current (quintessence-like or phantom) nature of DE, whether or not this can change in the past is determined by the value of $w_{a}$, given that at asymptotically early times ($z\to\infty$) the DE EoS tends to the value $w_{0}+w_{a}$. Finally, we note that models crossing which can cross between the two regimes are typically referred to as “quintom” models, and have been widely studied in the literature (Feng et al., 2006; Guo et al., 2005; Cai et al., 2007a, b; Zhang, 2009; Saridakis, 2010; Cai et al., 2012; Bahamonde et al., 2018; Leon et al., 2018; Panpanich et al., 2021), see Cai et al. (2010) for a review.
@@
$\frac{H^{2}(z)}{H_{0}^{2}}$ $=$ $\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}$ (3)
@@
where the density parameters satisfy $\Omega_{r}+\Omega_{m}+\Omega_{\Lambda}+\Omega_{x}=1$. Finally, as noted in Adil et al. (2023), it is natural to identify the combination of the evolving CPL component and the cosmological constant $\Lambda$ as comprising the combined DE sector, whose total density parameter is $\Omega_{\text{DE}}\equiv\Omega_{x}+\Omega_{\Lambda}$. The important thing to note is that, although $\Lambda$ itself and therefore $\Omega_{\Lambda}$ can be negative, the total DE density and therefore $\Omega_{\text{DE}}$ have to be positive in order to be able to drive cosmic acceleration and maintain agreement with cosmological observations. Roughly speaking, such a DE sector can in principle be compatible with cosmological observations provided $\Omega_{\text{DE}}\approx 0.7$, which of course is possible even if $\Omega_{\Lambda}<0$, as noted in several recent works (Dutta et al., 2020; Visinelli et al., 2019; Sen et al., 2022). We note that, in order to agree with observations, a more negative $\Omega_{\Lambda}$ needs to be compensated by more negative values of $w_{x}$, moving towards the phantom regime. These considerations lead to comparatively weak, order unity upper limits on $|\Omega_{\Lambda}|$, as discussed for instance in Visinelli et al. (2019) and Sen et al. (2022), whereas different combinations of $w_{x}$, $\Omega_{x}$, and $\Omega_{\Lambda}$ (the latter two summing to the same value of $\Omega_{\text{DE}}$) can lead to a very rich phenomenology, discussed in detail in Calderón et al. (2021) (see also Dash et al. (2024) for discussions on the expected sensitivity to a nCC from future 21-cm observations).
@@
From a theoretical standpoint, such a phenomenological model has been argued to loosely carry string motivation in the case where $\Lambda<0$ (Visinelli et al., 2019; Sen et al., 2022; Adil et al., 2023). While AdS vacua appear ubiquitously in string theory, the evolving DE component on top can, broadly speaking, be justified on the grounds that string compactifications typically predict the existence of a plethora of ultralight (pseudo)scalar particles. This usually goes under the name of “string axiverse” (see for example Svrcek & Witten, 2006; Arvanitaki et al., 2010; Cicoli et al., 2012; Visinelli & Vagnozzi, 2019; Cicoli et al., 2023), with the axion-like particles arising from Kaluza-Klein reduction of higher-dimensional form fields on the topological cycles of the compactification space. The topology of the compactification manifold fixes the number of particles, typically of order hundreds or more, and with masses spread over a huge number of decades, with the rough expectation that the distribution of the logarithms of the masses should be approximately uniform (Kamionkowski et al., 2014; Emami et al., 2016; Karwal & Kamionkowski, 2016), see also Marsh (2011); Ruchika et al. (2023) for related studies. The important thing to note is that the effective EoS of multiple interacting scalar fields can be phantom (Carroll et al., 2003; Vikman, 2005; Carroll et al., 2005; Deffayet et al., 2010; Sawicki & Vikman, 2013): this motivates the possibility of a component with positive energy density, but with EoS $w_{x}(z)$ potentially crossing $-1$ [as in Eq. (1) for suitable choices of $w_{0}$, $w_{a}$], sitting on top of a nCC.
[…]
@@
Before moving on we note that, once the combination of evolving CPL component and $\Lambda$ is identified as making up the combined DE sector, a natural quantity characterizing the latter is the effective equation of state $w_{\text{eff}}(z)$, which can be determined through the twice contracted Bianchi identity, and is given by (see Adil et al., 2023, for the full calculation):
{medsize}$\medsize w_{\rm eff}(z)=\frac{\Omega_{x}(1+z)^{2+3w_{0}+3w_{a}}\left[w_{0}+(w_{0}+w_{a})z\right]\exp\left(-3w_{a}\dfrac{z}{1+z}\right)-\Omega_{\Lambda}}{\Omega_{x}(1+z)^{3(1+w_{0}+w_{a})}\exp\left(-3w_{a}\frac{z}{1+z}\right)+\Omega_{\Lambda}}\,,\$
@@
for which in general $w_{\rm eff}(z=0)\neq w_{0}$ unless $\Omega_{\Lambda}=0$. It is also worth noting that if the values of $\Omega_{x}$, $\Omega_{\Lambda}$, $w_{0}$, and $w_{a}$ are such that the total DE energy density goes through zero and therefore switches sign at a certain redshift, the associated effective EoS given by Eq. (2) necessarily goes through a pole at the same redshift. This has been discussed in detail in Ozulker (2022); Adil et al. (2023), and does not in itself signal a pathology given that $w_{\rm eff}(z)$ is not associated to the dynamics of a single microscopical degree of freedom – however, this does highlight the importance of focusing on the total DE energy density rather than the effective EoS.
@@ 3 Methods
For each set of cosmological parameters discussed above, following Linder (2005) and as reported in Menci et al. (2020, 2022) and Adil et al. (2023), we compute the evolution of the matter density contrast $\delta$, from which we obtain the linear growth factor of density perturbations $D(z)$. We note that the evolution of $D(z)$ can differ significantly from that within $\Lambda$CDM. The reason is that the equation for the evolution of $\delta$ (see e.g. Adil et al., 2023) depends on both the normalized expansion rate $E(z)\equiv H(z)/H_{0}$ and its first derivative: a different background expansion therefore directly impacts the growth of structure, which within the model at hand can be either suppressed or enhanced depending on the choice of cosmological parameters, with the value of $\Omega_{\Lambda}$ playing an important role (see Adil et al., 2023, for more detailed discussions on this point).
@@
The next step is to compute the predicted maximal abundance of galaxies with stellar mass $M_{\star}$. More precisely, the relevant quantities are the comoving cumulative number or mass (stellar mass or DM halo mass) densities, which quantify the number or mass density of halos above a given threshold, or the number or stellar mass density of stars contained in galaxies more massive than a given threshold. Here we proceed as in Menci et al. (2020, 2022) and just recall the basic steps to compute these quantities, while encouraging the reader to consult the above papers for further details (see also Eqs. (3.7–3.10) of Adil et al., 2023, for explicit expressions of the four quantities discussed above). For a given set of the seven cosmological parameters we compute the DM halo mass function $dn(M,z)/dM$, which quantifies the number of DM haloes of mass $M$ per unit mass per unit comoving volume in the mass range $[M;M+dM]$ at a given redshift. We do so following the prescription of Sheth & Tormen (1999), itself an extension of the Press-Schechter formalism (Press & Schechter, 1974) accounting for ellipsoidal collapse.
@@
From the DM halo mass function we can obtain the maximal comoving number density of galaxies with stellar mass in the range $[M_{\star};M_{\star}+dM_{\star}]$, $dn_{\star}(M_{\star},z)/dM_{\star}$. This is given by $dn_{\star}(M_{\star},z)/dM_{\star}=f_{b}dn(M,z)/dM$, where $f_{b}\equiv\Omega_{b}/\Omega_{m}$ is the cosmic baryon fraction. We stress that this is the maximal density because it is computed under the extremely optimistic assumption that the entire available baryonic reservoir ends up being converted into stars. This assumption is of course unrealistic, as in reality the efficiency of converting gas into stars, typically denoted by $\epsilon$, is of order $\epsilon\lesssim 0.2$ or less, with a moderate redshift dependence (Leroy et al., 2008; Combes et al., 2011; Tacchella et al., 2018). However, this assumption is extremely conservative for the purposes of our study: given that galaxies cannot outnumber their DM haloes, one can exclude those cosmological models for which even when $\epsilon=1$ the predicted number density of galaxies of given stellar mass at a certain redshift falls short of the observed abundance of galaxies within the same range of redshift and stellar mass. To put it differently, a galaxy of stellar mass $M_{\star}$ can only form if a DM halo of mass $M_{\star}/\epsilon f_{b}$ has formed first. Finally, we can compute the corresponding cumulative comoving maximal number and stellar mass densities of galaxies with stellar masses larger than a given observational threshold $\overline{M_{\star}}$. These are obtained by integrating $dn(M,z)/dM$ and $f_{b}Mdn(M,z)/dM$ respectively, where the lower limit of the integration range is given by the threshold DM halo mass $\overline{M_{\star}}/f_{b}$ [^10] The upper limit of the integration range is formally $\infty$, although the integrand drops exponentially quickly for sufficiently large DM halo mass. and averaging over the cosmic volume $V(z_{1},z_{2})$ enclosed between the redshift range $z_{\min}-z_{\max}$ covered by the observations (the average is performed within the integral, again see Eqs. (3.7–3.10) of Adil et al., 2023, for explicit expressions of these quantities). Two further corrections are required before the computed cumulative comoving stellar mass density of galaxies, $\rho_{\star}(>\overline{M_{\star}},z)$, can be compared against the JWST observations, as we will discuss shortly.
@@
Our goal is now to compute $\rho_{\star}(>\overline{M_{\star}},z)$ as a function of various choices of cosmological parameters, while ensuring that the latter are in agreement with standard cosmological probes. This will allow us to discuss the complementarity between the latter, and the observed abundance of high-redshift galaxies from JWST. Operationally, we therefore proceed as follows. The cosmological datasets we consider are given below:  [^11] While more recent datasets are available for some of the measurements below (especially for what concerns BAO and SNeIa data), the reason why we chose these datasets was for consistency with the work of Sen et al. (2022), in order to make use of the chains which were produced therein. Nevertheless, we expect very minimal quantitative changes were we to use the most up-to-date eBOSS BAO data (Alam et al., 2021) and PantheonPlus SNeIa data (Scolnic et al., 2022), and no qualitative changes, given that our conclusions are almost entirely driven by the JWST observations. We therefore expect them to be robust against the use of slightly more update BAO and SNeIa data.
@@
  • CMB data: measurements of CMB temperature anisotropy and polarization power spectra, their cross-spectra, and CMB lensing power spectrum reconstructed from the temperature 4-point correlation function, from the Planck 2018 legacy release. This combination is usually referred to as TTTEEE+lowl+lowE+lensing, and we analyse it making use of the official Planck likelihood (Aghanim et al., 2020a, b, c).
@@
  • BAO data: isotropic and anisotropic distance and expansion rate measurements from the SDSS-MGS, 6dFGS, and BOSS DR12 collaborations. (Ross et al., 2015; Beutler et al., 2011; Alam et al., 2017)
@@
  • SNeIa data: distance moduli measurements from the Pantheon SNeIa sample within the redshift range $2.30.01<z<$ 2.3 (Scolnic et al., 2018).
  • SH0ES prior: a prior on the Hubble constant $H_{0}=(73.30\pm 1.04)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$, as determined by the SH0ES team (Riess et al., 2022).
@@
Considering the combination of the above datasets, we derive constraints on our set of cosmological parameters by making use of Markov Chain Monte Carlo (MCMC) methods, with predictions for the cosmological observables in question derived using the publicly available Boltzmann solver CLASS (Lesgourgues, 2011; Blas et al., 2011). We make use of the publicly available cosmological MCMC sampler MontePython 3.3 (Audren et al., 2013; Brinckmann & Lesgourgues, 2019). We monitor the convergence of the generated MCMC chains through the Gelman-Rubin parameter $R-1$ (Gelman & Rubin, 1992), requiring $R-1<0.02$ for our chains to be considered converged, and analyze them using the GetDist package (Lewis, 2019). For a full discussion of the methodology and the resulting constraints, we refer the reader to Sen et al. (2022).
@@
  • the six most massive, intrinsically red galaxies in the redshift range $9\lesssim z\lesssim 11$, identified in the first NIRCam observations of the JWST CEERS program, as reported in Labbé et al. (2023), and which we treat as a measurement of $\rho_{\text{obs}}(>\overline{M_{\star}})$ at $z_{\text{eff}}=10$;
  • the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts, in the redshift range $5\lesssim z\lesssim 6$, identified within the JWST FRESCO NIRCam/grism survey, as reported in Xiao et al. (2023), and which we treat as a measurement of $\rho_{\text{obs}}(>\overline{M_{\star}})$ at $z_{\text{eff}}=5.5$.
@@
From the MCMC chains described above, we then select only those points which are consistent within $2\sigma$ with the standard cosmological probes discussed earlier. We stress that in our MCMC analysis we have allowed for $\Omega_{\Lambda}\gtrless 0$, i.e. either a dS or AdS vacuum energy component. Each of the models consistent within $2\sigma$ with the standard cosmological observations is then compared against the inferred values of $\rho_{\text{obs}}(>\overline{M_{\star}})$ from the JWST CEERS and FRESCO observations described above. However, in order to properly carry out this comparison, we need to account (and correct) for the fact that both the values of $\rho_{\text{obs}}(>\overline{M_{\star}})$ reported by Labbé et al. (2023) and Xiao et al. (2023) have been obtained assuming a certain fiducial cosmology, in this case $\Lambda$CDM with a given choice of parameters. Let us refer to the vector of fiducial cosmology parameters as $\boldsymbol{f}$, whereas we denote by $\boldsymbol{\theta}$ the vector of cosmological parameters for each point in our MCMC chain. As discussed in detail in Menci et al. (2022), one then needs to apply two different corrections:
@@
We consider a given cosmological model (which, we recall, is already ensured to be consistent within $2\sigma$ with the standard cosmological probes) to be excluded if the maximal predicted stellar mass density of galaxies at the given effective redshift is lower than the value inferred from JWST CEERS or FRESCO observations. From this we can then compute the exclusion probability $P$ for a given model, or conversely the probability of consistency with the JWST observations $Q$ (obviously $P+Q=1$). This is performed using the procedure described in Menci et al. (2022), through a Monte Carlo procedure which also accounts for uncertainties in the observational estimates of masses (in particular potential systematics related to the spectral energy distribution fitting procedure), and statistical errors in the observational number densities. We refer the reader to Menci et al. (2022) for further details on the calculation of the exclusion probability.

### v2

@@ 2 Dark Energy models
Following the earlier work of Ref. [209], we consider a DE sector consisting of a cosmological constant $\Lambda\gtrless 0$ which in principle can take either sign (with positive or negative signs corresponding to a dS or AdS vacuum respectively), and with associated energy density $\rho_{\Lambda}=\Lambda/8\pi G$ of the same sign. On top of this dS or AdS vacuum we place an evolving DE component with strictly positive energy density $\rho_{x}(z)>0$. Rather than committing to a specific microphysical model for the evolving DE component, we assume that this is described by a time-evolving EoS $w_{x}(z)$ of the Chevallier-Polarski-Linder (CPL) form [210, 211]:
$w_{x}(z)=w_{0}+w_{a}\frac{z}{1+z}\,.$ (2.1)
There are various reasons why we adopt the widely used CPL parametrization, ranging from its highly manageable 2-dimensional nature, to its direct connection to several physical DE models (including several quintessence DE models, see e.g. Refs. [211, 212, 213, 214, 215]) and, last but not least, for ease of comparison to the earlier work of Ref. [209] [^8] See Ref. [216] for a recent discussion on potential shortcomings of the CPL parametrization. For completeness, we note that a number of other parametrizations for the EoS of dynamical DE components have been proposed in the literature, see e.g. Refs. [217, 218, 219, 220, 221, 222, 223, 224, 225, 226], see also Ref. [227] for discussions on the theoretical viability of these parametrizations. We refer to the evolving component being quintessence-like (phantom) at a given epoch if, at the redshift in question, $w(z)>-1$ ($<-1$). While $w_{0}$ controls the current (quintessence-like or phantom) nature of DE, whether or not this can change in the past is determined by the value of $w_{a}$, given that at asymptotically early times ($z\to\infty$) the DE EoS tends to the value $w_{0}+w_{a}$. Finally, we note that models crossing which can cross between the two regimes are typically referred to as “quintom” models, and have been widely studied in the literature [228, 229, 230, 231, 232, 233, 234, 235, 236, 237], see Ref. [238] for a review.
@@
$\frac{H^{2}(z)}{H_{0}^{2}}$ $=$ $\Omega_{r}(1+z)^{4}+\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}$ (2.3)
@@
where the density parameters satisfy $\Omega_{r}+\Omega_{m}+\Omega_{\Lambda}+\Omega_{x}=1$. Finally, as noted in Ref. [209], it is natural to identify the combination of the evolving CPL component and the cosmological constant $\Lambda$ as comprising the combined DE sector, whose total density parameter is $\Omega_{\text{DE}}\equiv\Omega_{x}+\Omega_{\Lambda}$. The important thing to note is that, although $\Lambda$ itself and therefore $\Omega_{\Lambda}$ can be negative, the total DE density and therefore $\Omega_{\text{DE}}$ have to be positive in order to be able to drive cosmic acceleration and maintain agreement with cosmological observations. Roughly speaking, such a DE sector can in principle be compatible with cosmological observations provided $\Omega_{\text{DE}}\approx 0.7$, which of course is possible even if $\Omega_{\Lambda}<0$, as noted in several recent works [190, 191, 195, 239, 240]. We note that, in order to agree with observations, a more negative $\Omega_{\Lambda}$ needs to be compensated by more negative values of $w_{x}$, moving towards the phantom regime. These considerations lead to comparatively weak, order unity upper limits on $|\Omega_{\Lambda}|$, as discussed for instance in Refs. [191, 195], whereas different combinations of $w_{x}$, $\Omega_{x}$, and $\Omega_{\Lambda}$ (the latter two summing to the same value of $\Omega_{\text{DE}}$) can lead to a very rich phenomenology, discussed in detail in Ref. [194] (see also Ref. [241] for discussions on the expected sensitivity to a nCC from future 21-cm observations).
@@
From a theoretical standpoint, such a phenomenological model has been argued to loosely carry string motivation in the case where $\Lambda<0$ [191, 195, 209]. While AdS vacua appear ubiquitously in string theory, the evolving DE component on top can, broadly speaking, be justified on the grounds that string compactifications typically predict the existence of a plethora of ultralight (pseudo)scalar particles. This usually goes under the name of “string axiverse” (see for example Refs. [242, 243, 244, 245, 246]), with the axion-like particles arising from Kaluza-Klein reduction of higher-dimensional form fields on the topological cycles of the compactification space. The topology of the compactification manifold fixes the number of particles, typically of order hundreds or more, and with masses spread over a huge number of decades, with the rough expectation that the distribution of the logarithms of the masses should be approximately uniform [247, 248, 249], see also Ref. [250, 192] for related studies. The important thing to note is that the effective EoS of multiple interacting scalar fields can be phantom [251, 161, 162, 252, 253]: this motivates the possibility of a component with positive energy density, but with EoS $w_{x}(z)$ potentially crossing $-1$ [as in Eq. (2.1) for suitable choices of $w_{0}$, $w_{a}$], sitting on top of a nCC.
Concerning the AdS vacuum itself, the exploratory results of Ref. [209] which our work seeks to confirm show that the relevant region of parameter space is one where $|\Omega_{\Lambda}|\sim{\cal O}(1)$, i.e. where the nCC is of the same order of the dS vacuum energy in $\Lambda$CDM (see also Refs. [190, 191, 195]). Therefore, the magnitude of the nCC is expected to be $\lesssim 10^{-123}$ in Planck units, and one could legitimately worry that this would introduce a (negative) cosmological constant problem. Intriguingly, recent work in string theory [254, 255] has led to the explicit construction of supersymmetric AdS4 vacua of the right magnitude. This has been achieved within the context of type IIB string theory in compactifications on orientifolds of Calabi-Yau threefold hypersurfaces, with the resulting solution preserving ${\cal N}=1$ supersymmetry, and with the key point of the construction being the perfect cancellation of all perturbative terms in the superpotential. Such a construction explicitly shows that it is possible to obtain an exponentially small nCC within string compactifications, thereby providing further string motivation for the region of parameter space we shall explore in this work.
@@
Before moving on we note that, once the combination of evolving CPL component and $\Lambda$ is identified as making up the combined DE sector, a natural quantity characterizing the latter is the effective equation of state $w_{\text{eff}}(z)$, which can be determined through the twice contracted Bianchi identity, and is given by (see Ref. [209] for the full calculation):
$w_{\rm eff}(z)=\frac{\Omega_{x}(1+z)^{2+3w_{0}+3w_{a}}\left[w_{0}+(w_{0}+w_{a})z\right]\exp\left(-3w_{a}\dfrac{z}{1+z}\right)-\Omega_{\Lambda}}{\Omega_{x}(1+z)^{3(1+w_{0}+w_{a})}\exp\left(-3w_{a}\frac{z}{1+z}\right)+\Omega_{\Lambda}}\,,\$
@@
for which in general $w_{\rm eff}(z=0)\neq w_{0}$ unless $\Omega_{\Lambda}=0$. It is also worth noting that if the values of $\Omega_{x}$, $\Omega_{\Lambda}$, $w_{0}$, and $w_{a}$ are such that the total DE energy density goes through zero and therefore switches sign at a certain redshift, the associated effective EoS given by Eq. (2) necessarily goes through a pole at the same redshift. This has been discussed in detail in Refs. [256, 209], and does not in itself signal a pathology given that $w_{\rm eff}(z)$ is not associated to the dynamics of a single microscopical degree of freedom – however, this does highlight the importance of focusing on the total DE energy density rather than the effective EoS.
@@ 3 Methods
For each set of cosmological parameters discussed above, following Ref. [257] and as reported in Refs. [88, 94, 209], we compute the evolution of the matter density contrast $\delta$, from which we obtain the linear growth factor of density perturbations $D(z)$. We note that the evolution of $D(z)$ can differ significantly from that within $\Lambda$CDM. The reason is that the equation for the evolution of $\delta$ (see e.g. Ref. [209]) depends on both the normalized expansion rate $E(z)\equiv H(z)/H_{0}$ and its first derivative: a different background expansion therefore directly impacts the growth of structure, which within the model at hand can be either suppressed or enhanced depending on the choice of cosmological parameters, with the value of $\Omega_{\Lambda}$ playing an important role (see Ref. [209] for more detailed discussions on this point).
@@
The next step is to compute the predicted maximal abundance of galaxies with stellar mass $M_{\star}$. More precisely, the relevant quantities are the comoving cumulative number or mass (stellar mass or DM halo mass) densities, which quantify the number or mass density of halos above a given threshold, or the number or stellar mass density of stars contained in galaxies more massive than a given threshold. Here we proceed as in Refs. [88, 94] and just recall the basic steps to compute these quantities, while encouraging the reader to consult the above papers for further details (see also see also Eqs. (3.7–3.10) of Ref. [209] for explicit expressions of the four quantities discussed above). For a given set of the seven cosmological parameters we compute the DM halo mass function $dn(M,z)/dM$, which quantifies the number of DM haloes of mass $M$ per unit mass per unit comoving volume in the mass range $[M;M+dM]$ at a given redshift. We do so following the prescription of Ref. [258], itself an extension of the Press-Schechter formalism [259] accounting for ellipsoidal collapse.
@@
From the DM halo mass function we can obtain the maximal comoving number density of galaxies with stellar mass in the range $[M_{\star};M_{\star}+dM_{\star}]$, $dn_{\star}(M_{\star},z)/dM_{\star}$. This is given by $dn_{\star}(M_{\star},z)/dM_{\star}=f_{b}dn(M,z)/dM$, where $f_{b}\equiv\Omega_{b}/\Omega_{m}$ is the cosmic baryon fraction. We stress that this is the maximal density because it is computed under the extremely optimistic assumption that the entire available baryonic reservoir ends up being converted into stars. This assumption is of course unrealistic, as in reality the efficiency of converting gas into stars, typically denoted by $\epsilon$, is of order $\epsilon\lesssim 0.2$ or less, with a moderate redshift dependence [260, 261, 262]. However, this assumption is extremely conservative for the purposes of our study: given that galaxies cannot outnumber their DM haloes, one can exclude those cosmological models for which even when $\epsilon=1$ the predicted number density of galaxies of given stellar mass at a certain redshift falls short of the observed abundance of galaxies within the same range of redshift and stellar mass. To put it differently, a galaxy of stellar mass $M_{\star}$ can only form if a DM halo of mass $M_{\star}/\epsilon f_{b}$ has formed first. Finally, we can compute the corresponding cumulative comoving maximal number and stellar mass densities of galaxies with stellar masses larger than a given observational threshold $\overline{M_{\star}}$. These are obtained by integrating $dn(M,z)/dM$ and $f_{b}Mdn(M,z)/dM$ respectively, where the lower limit of the integration range is given by the threshold DM halo mass $\overline{M_{\star}}/f_{b}$ [^10] The upper limit of the integration range is formally $\infty$, although the integrand drops exponentially quickly for sufficiently large DM halo mass. and averaging over the cosmic volume $V(z_{1},z_{2})$ enclosed between the redshift range $z_{\min}-z_{\max}$ covered by the observations (the average is performed within the integral, again see Eqs. (3.7–3.10) of Ref. [209] for explicit expressions of these quantities). Two further corrections are required before the computed cumulative comoving stellar mass density of galaxies, $\rho_{\star}(>\overline{M_{\star}},z)$, can be compared against the JWST observations, as we will discuss shortly.
@@
Our goal is now to compute $\rho_{\star}(>\overline{M_{\star}},z)$ as a function of various choices of cosmological parameters, while ensuring that the latter are in agreement with standard cosmological probes. This will allow us to discuss the complementarity between the latter, and the observed abundance of high-redshift galaxies from JWST. Operationally, we therefore proceed as follows. The cosmological datasets we consider are given below:  [^11] While more recent datasets are available for some of the measurements below (especially for what concerns BAO and SNeIa data), the reason why we chose these datasets was for consistency with the work of Ref. [195], in order to make use of the chains which were produced therein. Nevertheless, we expect very minimal quantitative changes were we to use the most up-to-date eBOSS BAO data [263] and PantheonPlus SNeIa data [264], and no qualitative changes, given that our conclusions are almost entirely driven by the JWST observations. We therefore expect them to be robust against the use of slightly more update BAO and SNeIa data.
@@
  • CMB data: measurements of CMB temperature anisotropy and polarization power spectra, their cross-spectra, and CMB lensing power spectrum reconstructed from the temperature 4-point correlation function, from the Planck 2018 legacy release. This combination is usually referred to as TTTEEE+lowl+lowE+lensing, and we analyse it making use of the official Planck likelihood [265, 266, 267].
@@
  • BAO data: isotropic and anisotropic distance and expansion rate measurements from the SDSS-MGS, 6dFGS, and BOSS DR12 collaborations [268, 269, 270].
@@
  • SNeIa data: distance moduli measurements from the Pantheon SNeIa sample within the redshift range $2.30.01<z<$ 2.3 [271].
  • SH0ES prior: a prior on the Hubble constant $H_{0}=(73.30\pm 1.04)\,{\text{km}}/{\text{s}}/{\text{Mpc}}$, as determined by the SH0ES team [272].
@@
Considering the combination of the above datasets, we derive constraints on our set of cosmological parameters by making use of Markov Chain Monte Carlo (MCMC) methods, with predictions for the cosmological observables in question derived using the publicly available Boltzmann solver CLASS [273, 274]. We make use of the publicly available cosmological MCMC sampler MontePython 3.3 [275, 276]. We monitor the convergence of the generated MCMC chains through the Gelman-Rubin parameter $R-1$ [277], requiring $R-1<0.02$ for our chains to be considered converged, and analyze them using the GetDist package [278]. For a full discussion of the methodology and the resulting constraints, we refer the reader to Ref. [195].
@@
  • the six most massive, intrinsically red galaxies in the redshift range $9\lesssim z\lesssim 11$, identified in the first NIRCam observations of the JWST CEERS program, as reported in Ref. [28], and which we treat as a measurement of $\rho_{\text{obs}}(>\overline{M_{\star}})$ at $z_{\text{eff}}=10$;
  • the three most massive, optically dark (dust-obscured) galaxies with robust spectroscopic redshifts, in the redshift range $5\lesssim z\lesssim 6$, identified within the JWST FRESCO NIRCam/grism survey, as reported in Ref. [29], and which we treat as a measurement of $\rho_{\text{obs}}(>\overline{M_{\star}})$ at $z_{\text{eff}}=5.5$.
@@
From the MCMC chains described above, we then select only those points which are consistent within $2\sigma$ with the standard cosmological probes discussed earlier. We stress that in our MCMC analysis we have allowed for $\Omega_{\Lambda}\gtrless 0$, i.e. either a dS or AdS vacuum energy component. Each of the models consistent within $2\sigma$ with the standard cosmological observations is then compared against the inferred values of $\rho_{\text{obs}}(>\overline{M_{\star}})$ from the JWST CEERS and FRESCO observations described above. However, in order to properly carry out this comparison, we need to account (and correct) for the fact that both the values of $\rho_{\text{obs}}(>\overline{M_{\star}})$ reported by Ref. [28] and Ref. [29] have been obtained assuming a certain fiducial cosmology, in this case $\Lambda$CDM with a given choice of parameters. Let us refer to the vector of fiducial cosmology parameters as $\boldsymbol{f}$, whereas we denote by $\boldsymbol{\theta}$ the vector of cosmological parameters for each point in our MCMC chain. As discussed in detail in Ref. [94], one then needs to apply two different corrections:
@@
We consider a given cosmological model (which, we recall, is already ensured to be consistent within $2\sigma$ with the standard cosmological probes) to be excluded if the maximal predicted stellar mass density of galaxies at the given effective redshift is lower than the value inferred from JWST CEERS or FRESCO observations. From this we can then compute the exclusion probability $P$ for a given model, or conversely the probability of consistency with the JWST observations $Q$ (obviously $P+Q=1$). This is performed using the procedure described in Ref. [94], through a Monte Carlo procedure which also accounts for uncertainties in the observational estimates of masses – in particular potential systematics related to the spectral energy distribution (SED) fitting procedure – and statistical errors in the observational number densities.
[…]